# A Liquid-Nitrogen-Cooled ${}^{40}\mathrm{Ca}^{+}$ Ion Optical Clock with a Systematic Uncertainty of $4.4\times 10^{-19}$

### v2

@@ I ACKNOWLEDGMENTS
We are deeply grateful to Dr. XXX for his valuable discussions, and to Professor XXX for their careful review of the manuscript. This work is supported by the National Key R&D Program of China (Grant No. 2022YFB3904001), Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0300901), the National Natural Science Foundation of China (Grant Nos. 12204494, 12121004, 12320101003 and 11934014), the CAS Youth Innovation Promotion Association (Grant No Y2022099), the Natural Science Foundation of Hubei Province (Grant Nos. 2022CFA013 and 2023EHA006), and the CAS Project for Young Scientists in Basic Research (Grant Nos. YSBR-055 and YSBR-085).

### v4

@@ Abstract
Acknowledgements—We are deeply grateful to Bing Yan for providing the Ca+–He potential data, and to A. C. Vutha for helpful discussions. This work is supported by the National Key R&D Program of China (Grant Nos. 2022YFB3904001 and 2022YFB3904004), Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0300901), the National Natural Science Foundation of China (Grant Nos. 12204494, 12121004, 12320101003, 11934014, 12374235 and 12174402), the CAS Youth Innovation Promotion Association (Grant No Y2022099), the Natural Science Foundation of Hubei Province (Grant Nos. 2022CFA013 and 2023EHA006), the CAS Project for Young Scientists in Basic Research (Grant Nos. YSBR-055 and YSBR-085), and Strategic Priority Research Program of the Chinese Academy of Sciences (XDB0920101 and XDB0920202).